# Universal logical operations with a dynamical qubit in Floquet code

### v2

@@ Abstract
Quantum error correction (QEC) protects quantum systems against inevitable noises and control inaccuracies, providing a pathway towards fault-tolerant (FT) quantum computation. However, the significant overhead of physical qubits required to encode a single logical qubit poses a major challenge for scalability and practical implementation. Floquet QEC codes, a recent innovation, mitigate this challenge by utilizing time-periodic measurements to introduce additional dynamical logical qubits, thereby enhancing resource efficiency in QEC. Here, we experimentally implement the Floquet-Bacon-Shor code on a superconducting quantum processor. We encode a dynamical logical qubit within a $3\times 3$ lattice of data qubits, alongside a conventional static logical qubit. We demonstrate FT encoding and measurement of the two-qubit logical states and stabilize the states using repeated error detection cycles. Additionally, we showcase universal single-qubit logical gates on the dynamical qubit. By implementing a logical CNOT gate, we entangle the dynamical and static logical qubits, generating an error-detected logical Bell state with a fidelity of 75.9%. Our results highlight the potential of Floquet codes for scalable and resource-efficient FT quantum computation.
I Introduction
@@
Figure 1: Implementation of the FBS code. a, Schematic of the quantum device for the FBS code, consisting of 9 data qubits (white circles labeled as $D_{i}$, where i = 1 to 9) and 12 ancilla qubits (orange circles for $Z$ ancillas and blue circles for $X$ ancillas). The four central qubits are alternately used as $X$ and $Z$ ancillas in different stabilizer rounds. Extra ancilla qubits (grey circles) are used for circuit compression and the construction of logical CNOT gate. Each pair of qubits is connected by a tunable coupler (H shape). b, Illustration of stabilizer measurements for the BS code (left) and the FBS code (right). The FBS code features period-4 stabilizer measurements, during which the logical operators $X_{\rm d}$ (purple lines) and $Z_{\rm d}$ (green lines) for the dynamical qubit evolve across different rounds. In contrast, the logical operators $X_{\rm s}$ and $Z_{\rm s}$ for the static qubit remain fixed and are identical in both codes. Weight-2 Pauli measurements are used to check the parity of the corresponding data qubits in the $X$ (blue) and $Z$ (orange) bases. c, Cumulative distributions of simultaneous errors for 21 single-qubit gates, 32 CZ gates, ancilla qubit readout, and data qubit idle (with dynamical decoupling) during ancilla readout are depicted. The median error rates are 0.07%, 0.97%, 1.58%, and 1.43%, respectively, as indicated by the vertical lines.
@@
Quantum computers hold potential for solving computational problems that are provably intractable for classical systems, with transformative applications in quantum simulation [1], cryptography [2], and machine learning [3]. However, their practical realization is hindered by a fundamental challenge: quantum information is inherently fragile, susceptible to inevitable environmental noise and control inaccuracies. To address this challenge, quantum error correction (QEC) has been established as a critical framework [4, 5, 6]. By redundantly encoding logical qubits into larger systems with many physical qubits, QEC detects and corrects errors through repeated syndrome measurements. Combined with fault-tolerant (FT) logical operations, QEC not only safeguards quantum information, but also lays the foundation for scalable and reliable quantum computation.
A variety of QEC codes have been proposed, including surface code [7], color code [8], Bacon-Shor (BS) code [9], and quantum LDPC codes [10, 11]. Significant experimental progresses have been achieved in recent years across various platforms [12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22]. Notably, exponential error suppression has been demonstrated by scaling up a surface-code logical qubit using superconducting circuits [23]. While high-performance logical qubits can be realized by increasing the size of quantum processors, the code rate–the ratio of logical qubits to physical qubits–remains a crucial metric [11]. Low encoding efficiency can make FT quantum computing resource-intensive, especially for solid-state quantum devices limited to nearest-neighbor coupling.
Recently proposed time-dynamical QEC, including Floquet codes [24, 25] and generalized time-dynamical code implementations [26, 27, 28, 29, 30], opens up new opportunities for fault-tolerant quantum computation. A recent work [31] has shown the flexibility in surface code realization using time-dynamical approaches. Unlike conventional stabilizer or subsystem codes that encode logical information in a fixed subspace, the Floquet codes [32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52] allow logical information to evolve periodically over time. Additional dynamical logical qubits are encoded through an appropriate measurement schedule, thereby improving the code rate. During the periodic schedule of Pauli measurements, logical subspaces and operators adapt to stabilizer measurements while preserving the encoded information.
@@
In this work, we implement the Floquet-Bacon-Shor (FBS) code [44] on a superconducting quantum processor, encoding one dynamical logical qubit and one conventional static logical qubit simultaneously within a $3\times 3$ array of data qubits. We demonstrate FT encoding and measurement of the two-qubit logical states, and stabilize these states using multi-round stabilizer measurements. We showcase FT Pauli gates and non-fault-tolerant (nFT) arbitrary rotation gates on the dynamical qubit. Furthermore, we achieve a logical Bell state with a fidelity of $75.9\%$ by entangling the dynamical and static logical qubits via a logical CNOT gate, with the gate fidelity assessed at $84.1\%$ through logical quantum process tomography (LQPT). With these demonstrated universal two-qubit logical operations, our results highlight the potential of Floquet codes for scalable and resource-efficient QEC.
II Results
@@
Figure 2: Encoding, stabilization, and measurement of two-qubit logical states in the FBS code. a, Simplified circuit for encoding, stabilizing and measuring two-qubit logical states. All physical qubits are initialized to the ground state before the circuit. Period-4 stabilizer measurements are alternately performed following the encoding circuit. b, Encoding circuits for logical Pauli states $|+,0\rangle$ and $|1,0\rangle$. BS stabilizer measurements are included in the latter case, with the dynamical qubit state post-selected to be $|0\rangle$. c, Example logical measurement circuits after stabilizer $S_{Z}^{\rm D}$: FT $ZZ$ measurement (left) and nFT $XZ$ measurement (right). Logical operators are indicated within the dashed-line boxes. Measurement outcomes of some data qubits are used to calculate the values of the $X$ (blue) and $Z$ (orange) stabilizers, enabling error detection. The notation $D_{147}$ denotes the three data qubits $D_{1}$, $D_{4}$, and $D_{7}$, and so on. d, LQST fidelities of all 36 two-qubit Pauli states. The 16 states $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}^{\otimes 2}$ are fault-tolerantly prepared and exhibit higher fidelity, as shown on the left side (white background). The blue portions indicate the fidelities without error detection, while the green portions highlight the fidelity improvement with error detection in the logical measurement. e, Experimentally measured (filled symbols) and simulated (open symbols) expectation values of $X_{\rm s}Z_{\rm d}$ operator versus stabilizer round, for the encoded states $|-,1\rangle$ (top) and $|-,0\rangle$ (bottom). The solid and dashed lines represent exponential fits. Each stabilizer round lasts 920 ns, comprising 720 ns for readout pulses and resonator photon depletion, and 200 ns for single-qubit and two-qubit gates. The cyan and green colors correspond to data without post-processing (raw) and data with measured error syndromes post-selected out (detection), respectively. Due to the exponential decrease in the retained data during error detection, the analysis is limited to 16 rounds, with a total of 100,000 experimental shots. The insets show the error detection probabilities for the weight-6 stabilizers over 24 measurement rounds, with the final round calculated from the logical measurements.
@@
The FBS code we demonstrate is a generalization of the $[[9,1,3]]$ BS code [9], a subsystem code that encodes one logical qubit on a $3\times 3$ square lattice, as illustrated in Fig. 1a. Each vertex represents a data qubit, and each edge corresponds to a weight-2 Pauli measurement. The static logical qubit is encoded in a fixed subsystem, with its logical Pauli operators given by
@@
The BS code can be stabilized in a single measurement round using only weight-2 Pauli measurements, whose ordering is inconsequential [9], as shown in Fig. 1b. The four weight-6 stabilizers, $S_{X}^{\text{A}}$, $S_{Z}^{\text{B}}$, $S_{X}^{\text{C}}$, and $S_{Z}^{\text{D}}$, are derived by multiplying the outcomes of the corresponding Pauli measurements, enabling the detection and subsequent correction of any single-qubit error. Since these Pauli measurements do not always commute, the Hilbert space of the 9 physical qubits consists of 4 degrees of freedom fixed by the stabilizers, 1 degree for logical information, and 4 additional degrees corresponding to the gauge space, which does not encode logical information. By imposing a period-4 ordering on the weight-2 Pauli measurements, the gauge space dynamically encodes an additional logical qubit, thereby forming the FBS code. In each round, $4$ weight-$2$ Pauli measurements are performed, whose outcomes determine one weight-$6$ stabilizer and also influence the evaluation of dynamical operators (see Methods). The logical Pauli operators $X_{\text{d}}$ and $Z_{\text{d}}$ of the dynamically encoded qubit evolve periodically during the stabilizer measurements, as indicated by the purple and green lines in Fig. 1b. Throughout this paper, the initial operator definitions are consistent with stabilizer round A. Consequently, the first stabilizer round A does not change the encoded operators. Both the static and dynamical logical qubits are stabilized over the 4 measurement rounds. Although the $3\times 3$ system size is insufficient for error correction of the dynamical qubit, the stabilizers extracted from the measurements are sufficient to detect any single-qubit error. Within the framework of error detection, an operation is considered FT if any single-qubit error produces a non-trivial syndrome and is thereby detectable [53].
@@
We implement the FBS code on a superconducting quantum processor comprising a $6\times 11$ array of Transmon qubits [54], each coupled to its four nearest neighbors through tunable couplers [55]. The experimental setup and device parameters are detailed in Supplementary Information Section II. The static and dynamical logical qubits are encoded on a 21-qubit subset of the processor, consisting of $N_{d}=9$ data qubits and $N_{a}=12$ ancilla qubits. The data qubits, labeled as $D_{i}$ ($i=1$ to 9), form a $3\times 3$ array and are interlaced with $X$ and $Z$ ancilla qubits for stabilizer measurements. Additional ancilla qubits are incorporated to optimize the QEC circuits. High-fidelity quantum gates and quantum non-demolition (QND) readout [56] are implemented on the 21-qubit subset, with the cumulative distributions of errors illustrated in Fig. 1c. The readout fidelities of the ancilla qubits are benchmarked using repeated measurements [57] (see Supplementary Information Section III). Dynamical decoupling (DD) techniques [58] are employed to protect the data qubits from low-frequency noise and cross-measurement-induced dephasing [59, 60, 61]. The idle errors (with DD) on the data qubits during ancilla qubit readout are quantified using interleaved randomized benchmarking [62].
@@
Logical measurements are performed by measuring the data qubits in their appropriate bases, as illustrated by the example circuits in Fig. 2c. The measurement circuits for the dynamical qubit change according to the logical operator definitions in different stabilizer rounds. The outcome of a logical operator is determined by calculating the total parity of the relevant data qubits (see Methods). Data qubits that do not define the logical operators are also measured to enable error detection. The $Z_{\rm s}Z_{\rm d}$ measurement is FT because the measured outcomes allow for the calculation of stabilizers $S_{Z}^{\rm D}$ and $S_{Z}^{\rm B}$, enabling the detection of any single-qubit error. In contrast, the measurement of $X_{\rm s}Z_{\rm d}$ is nFT; however, the value of stabilizer $S_{X}^{\rm C}$ can be calculated from the measurement outcomes to detect phase-flip errors and measurement errors on qubits $D_{1}$ through $D_{6}$. We present measured fidelities of the 36 encoded states in Fig. 2d. The state fidelities are determined using logical quantum state tomography (LQST), where 9 logical bases $\{Z,X,Y\}^{\otimes 2}$ are measured (see Methods for details on LQST and Supplementary Information Section VI for full measurement circuits). Error detection during the logical measurements improves the state fidelities, as indicated by the green portions in the plot.
Figure 3: Logical gates on the dynamical qubit. a, FT Pauli gates on the dynamical qubit. Transversal Pauli gates are applied to the data qubits and inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, as shown at the top. The final state is characterized using LQST. The state infidelities with error detection after applying the Pauli gates on different encoded states are shown below the circuit. For each dynamical qubit state, the infidelity is averaged over 4 static qubit states {$|0\rangle$, $|1\rangle$, |+ket|$\rangle$| + ⟩, |ket|$\rangle$| - ⟩}. Error bars correspond to 95% confidence intervals. b, Logical rotation gates on the dynamical qubit around the $Z_{\rm d}$-axis (left) and $X_{\rm d}$-axis (right). The measurement expectation values of the operators $X_{\rm d}$ (cyan), $Y_{\rm d}$ (green) and $Z_{\rm d}$ (orange) are shown as functions of rotation angles $\varphi$ and $\theta$. The gate circuits (shown at the top) for rotations around the $Z_{\rm d}$-axis ($X_{\rm d}$-axis) are inserted between stabilizers $S_{Z}^{\rm B}$ ($S_{X}^{\rm C}$) and $S_{X}^{\rm C}$ ($S_{Z}^{\rm D}$). The filled (open) symbols represent data with (without) error detection in the four stabilizer rounds and the final logical measurement. The dashed curves correspond to trigonometric fits.
@@
We perform period-4 stabilizer measurements to preserve the encoded two-qubit logical states. Specifically, the four stabilizers $S_{X}^{\rm A}$, $S_{Z}^{\rm B}$, $S_{X}^{\rm C}$, and $S_{Z}^{\rm D}$ are alternately performed to detect bit-flip or phase-flip errors in both the static and dynamical qubits. Each measurement result is compared to the result of the same stabilizer type from four rounds earlier, with the first four rounds compared to the encoding value. Errors during this process manifest as changes in the measurement outcomes, and the corresponding data are post-selected out in error detection. We present the preservation results for the encoded states $|-,0\rangle$ and $|-,1\rangle$ in Fig. 2e, showing the measured expectation values of the operator $X_{\rm s}Z_{\rm d}$. The logical error rates ($\varepsilon_{\rm exp}$) for these states, extracted by fitting the exponential decay of the expectation values, are reduced from 14.7% and 14.3% to 2.6% and 2.5%, respectively. This improvement corresponds to an increase in the logical lifetime $T_{\rm L}=\tau_{\rm exp}/2\varepsilon_{\rm exp}$ from 3.1 $\mu$s and 3.2 $\mu$s to 17.7 $\mu$s and 18.4 $\mu$[17], respectively. Here, $\tau_{\rm exp}=920$ ns is the duration of a stabilizer round. Numerical simulation results using the open-source package Stim [65], based on monitored gate and readout errors, are also shown for comparison. The larger error rates observed in experiments may arise from leakage-induced correlated errors [66, 67, 68, 60, 69, 70, 64]. We measure an average leakage rate of 0.18% per round on the physical qubits using three-state readout at the end of the circuit (see Supplementary Information IV). In both the experimental and simulation data, round-dependent oscillations are observed. Specifically, higher expectation values occur every four rounds, corresponding to the completion of a full error detection cycle. Beyond error detection, error correction can be applied to the static qubit alone (see Supplementary Information VI). However, the effectiveness of error correction diminishes for the two-qubit states due to the small code distance of the dynamical qubit and the correlation between errors in the two logical qubits.
We demonstrate universal logical gates on the dynamical qubit, including FT Pauli gates and nFT arbitrary rotation gates. For logical gates on the static qubit, see Supplementary Information Section V and reference [12]. The Pauli gates are implemented transversally using parallel single-qubit gates on the corresponding data qubits, as defined by the logical operators. We insert the Pauli gates {$I_{\rm s}\otimes X_{\rm d}$, $I_{\rm s}\otimes Y_{\rm d}$, $I_{\rm s}\otimes Z_{\rm d}$} between the stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, as illustrated in Fig 3a. The final states, after applying the gates and conducting four stabilizer rounds, are characterized using LQST. For each encoded Pauli state of the dynamical qubit, the static qubit traverses four states: $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}$. The average state infidelities after applying the gates $I_{\rm s}\otimes X_{\rm d}$, $I_{\rm s}\otimes Y_{\rm d}$, and $I_{\rm s}\otimes Z_{\rm d}$ are $0.255\pm 0.011$, $0.263\pm 0.013$, and $0.256\pm 0.011$, respectively. These values are close to the infidelity $0.253\pm 0.011$ observed for the identity operation (i.e., no gates inserted), indicating that the FT Pauli gates induce negligible errors.
The construction of arbitrary rotation gates is facilitated by the use of ancilla qubits (see Supplementary Information Section I for theoretical details). For logical rotations around the $Z_{\rm d}$-axis ($R_{Z}^{\varphi}$), a physical gate $R_{z}^{\varphi}$ is inserted on an ancilla qubit and placed between four physical CNOT gates acting on $D_{7}$ and $D_{8}$, as illustrated at the top of Fig. 3b. The entire circuit is then inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$. Similarly, logical rotations around the $X_{\rm d}$-axis ($R_{X}^{\theta}$) are implemented using an $R_{z}^{\theta}$ gate on an ancilla qubit, and the corresponding circuit is inserted between stabilizers $S_{X}^{\rm C}$ and $S_{Z}^{\rm D}$. We sweep the rotation angles $\varphi$ and $\theta$ and measure the dynamical qubit in three bases: $X_{\rm d}$, $Y_{\rm d}$, and $Z_{\rm d}$. The measurement results are shown in Fig. 3b, with filled (open) symbols representing data with (without) error detection. With error detection enabled, the lower amplitude in the $Y_{\rm d}$-direction is primarily due to the nFT nature of the logical $Y_{\rm d}$ measurement. Without error detection, the amplitude in the $Y_{\rm d}$-direction is lower than that in the $Z_{\rm d}$-direction for the $R_{X}^{\theta}$ rotation, due to more phase-flip errors propagating through the nFT $R_{X}^{\theta}$ gate.

### v3

# Logical operations with a dynamical qubit in Floquet-Bacon-Shor code

@@ Abstract
Quantum error correction (QEC) protects quantum systems against inevitable noises and control inaccuracies, providing a pathway towards fault-tolerant (FT) quantum computation. Stabilizer codes, including surface code and color code, have long been the focus of research and have seen significant experimental progress in recent years. Recently proposed time-dynamical QEC, including Floquet codes and generalized time-dynamical code implementations, opens up new opportunities for FT quantum computation. By employing a periodic schedule of low-weight parity checks, Floquet codes can generate additional dynamical logical qubits, offering enhanced error correction capabilities and potentially higher code performance. Here, we experimentally implement the Floquet-Bacon-Shor code on a superconducting quantum processor. We encode a dynamical logical qubit within a $3\times 3$ lattice of data qubits, alongside a conventional static logical qubit. We demonstrate FT encoding and measurement of the two-qubit logical states, and stabilize these states using repeated error detection. We showcase universal single-qubit logical gates on the dynamical qubit. Furthermore, by implementing a logical CNOT gate, we entangle the dynamical and static logical qubits, generating an error-detected logical Bell state with a fidelity of 75.9%. Our results highlight the potential of Floquet codes for resource-efficient FT quantum computation.
@@
Figure 1: Implementation of the FBS code. (a) Schematic of the quantum device for the FBS code, consisting of 9 data qubits (white circles labeled as $D_{i}$, where i = 1 to 9) and 12 ancilla qubits (orange circles for $Z$ ancillas and blue circles for $X$ ancillas). The four central qubits are alternately used as $X$ and $Z$ ancillas in different stabilizer rounds. Extra ancilla qubits (grey circles) are used for circuit compression and the construction of logical CNOT gate. Each pair of qubits is connected by a tunable coupler (H shape). (b) Illustration of stabilizer measurements for the BS code (left) and the FBS code (right). The FBS code features period-4 stabilizer measurements, during which the logical operators $X_{\rm d}$ (purple lines) and $Z_{\rm d}$ (green lines) for the dynamical qubit evolve across different rounds. In contrast, the logical operators $X_{\rm s}$ and $Z_{\rm s}$ for the static qubit remain fixed and are identical in both codes. Weight-2 Pauli measurements are used to check the parity of the corresponding data qubits in the $X$ (blue) and $Z$ (orange) bases. (c) Cumulative distributions of simultaneous errors for single-qubit gates, CZ gates, ancilla qubit readout, and data qubit idle (with dynamical decoupling) during ancilla readout are depicted. The median error rates are 0.07%, 0.97%, 1.58%, and 1.43%, respectively, as indicated by the vertical lines.
@@
Quantum computers hold potential for solving computational problems that are provably intractable for classical systems [1, 2, 3]. However, their practical realization is hindered by a fundamental challenge: quantum information is inherently fragile, susceptible to inevitable environmental noise and control inaccuracies. To address this challenge, quantum error correction (QEC) has been established as a critical framework [4, 5, 6]. By redundantly encoding logical qubits into larger systems with many physical qubits, QEC detects and corrects errors through repeated stabilizer measurements. QEC codes, including surface code [7], color code [8], Bacon-Shor (BS) code [9], and quantum LDPC code [10], have made significant experimental progress across various systems in recent years [11, 12, 13, 14, 15, 16, 17, 18, 19, 20]. Particularly in logical state preservation, exponential error suppression has been achieved by scaling a surface-code logical qubit in superconducting circuits [21]. Beyond state preservation, logical operations—such as arbitrary single-qubit gates and entangling gates between logical qubits—are essential for fault-tolerant (FT) quantum computing [22]. However, implementing these operations is generally challenging, especially in physical systems with local connectivity [23, 24, 25].
Recently proposed time-dynamical QEC, including Floquet codes [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49] and generalized time-dynamical code implementations [50, 51, 52, 53, 54, 55], opens up new opportunities for FT quantum computation by transcending the usual paradigm of stabilizer codes or subsystem codes, where logical information is encoded in a fixed subspace. The Floquet codes hold the advantage of requiring only low-weight Pauli measurements, typically of weight-2 [26], which are naturally compatible with physical architectures supporting native two-qubit measurements [56, 57]. Moreover, by periodically encoding logical information over time, Floquet codes enable additional dynamical logical qubits compared to their parent codes, offering greater flexibility and potentially higher code performance [39].
@@
In this work, we implement the Floquet-Bacon-Shor (FBS) code [41] on a superconducting quantum processor, encoding one dynamical logical qubit and one conventional static logical qubit simultaneously within a $3\times 3$ array of data qubits. We demonstrate FT encoding and measurement of the two-qubit logical states, and stabilize these states using multi-round stabilizer measurements. We showcase FT Pauli gates and non-fault-tolerant (nFT) arbitrary rotation gates on the dynamical qubit. Furthermore, we achieve a logical Bell state with a fidelity of $75.9\%$ by entangling the dynamical and static logical qubits via a logical CNOT gate, with the gate fidelity assessed at $84.1\%$ through logical quantum process tomography (LQPT). Our results highlight the potential of Floquet codes for scalable and resource-efficient QEC.
@@
Figure 2: Encoding, stabilization, and measurement of two-qubit logical states in the FBS code. (a) Simplified circuit for encoding, stabilizing and measuring two-qubit logical states. All physical qubits are initialized to the ground state before the circuit. Period-4 stabilizer measurements are alternately performed following the encoding circuit. (b) Encoding circuits for logical Pauli states $|+,0\rangle$ and $|1,0\rangle$. BS stabilizer measurements are included in the latter case, with the dynamical qubit state post-selected to be $|0\rangle$. The notation $D_{147}$ denotes the three data qubits $D_{1}$, $D_{4}$, and $D_{7}$, and so on. (c) Example logical measurement circuits after stabilizer $S_{Z}^{\rm D}$: FT $ZZ$ measurement (left) and nFT $XZ$ measurement (right). Logical operators are indicated within the dashed-line boxes. Measurement outcomes of data qubits are used to calculate the values of the $X$ (blue) and $Z$ (orange) stabilizers, enabling error detection. (d) LQST fidelities of all 36 two-qubit Pauli states. The 16 states $\{|0\rangle,|1\rangle,|+\rangle,|-\rangle\}^{\otimes 2}$ are fault-tolerantly prepared and exhibit higher fidelity, as shown on the left side (white background). The blue portions indicate the raw fidelities (i.e., without using syndrome data), while the green portions highlight the fidelity improvement by error detection (i.e., with syndrome data post-selected). (e) Experimentally measured (filled symbols) and simulated (open symbols) expectation values of $X_{\rm s}Z_{\rm d}$ operator versus stabilizer round, for the encoded states $|-,1\rangle$ (top) and $|-,0\rangle$ (bottom). The solid and dashed lines represent exponential fits. Each stabilizer round lasts 920 ns, comprising 720 ns for ancilla readout and 200 ns for single-qubit and two-qubit gates. The cyan and green colors correspond to the raw data and error-corrected data, respectively. The dashed grey lines indicate the average error rates of physical two-qubit states. Due to the exponential decrease in the retained data during error detection, the analysis is limited to 16 rounds, with a total of 100,000 experimental shots. The insets show the error detection probabilities for the weight-6 stabilizers over 24 measurement rounds, with the final round calculated from the logical measurements.
@@
The FBS code we demonstrate is a generalization of the $[[9,1,3]]$ BS code [9], a subsystem code that encodes one logical qubit on a $3\times 3$ square lattice, as illustrated in Fig. 1(a). Each vertex represents a data qubit, and each edge corresponds to a weight-2 Pauli measurement. The static logical qubit is encoded in a fixed subsystem, with its logical Pauli operators given by
@@
The BS code can be stabilized in a single measurement round using only weight-2 Pauli measurements, whose ordering is inconsequential [9], as shown in Fig. 1(b). The four weight-6 stabilizers, $S_{X}^{\text{A}}$, $S_{Z}^{\text{B}}$, $S_{X}^{\text{C}}$, and $S_{Z}^{\text{D}}$, are derived by multiplying the outcomes of the corresponding Pauli measurements, enabling the detection and subsequent correction of any single-qubit error. Since these Pauli measurements do not always commute, the Hilbert space of the 9 physical qubits consists of 4 degrees of freedom fixed by the stabilizers, 1 degree for logical information, and 4 additional degrees corresponding to the gauge space, which does not encode logical information. By imposing a period-4 ordering on the weight-2 Pauli measurements, the gauge space dynamically encodes an additional logical qubit, thereby forming the FBS code. The logical Pauli operators $X_{\text{d}}$ and $Z_{\text{d}}$ of the dynamically encoded qubit evolve periodically during the stabilizer measurements, as indicated by the purple and green lines in Fig. 1(b). Throughout this paper, the initial operator definitions are consistent with stabilizer round A. Consequently, the first stabilizer round A does not change the encoded operators. Both the static and dynamical logical qubits are stabilized over the 4 measurement rounds. Although the $3\times 3$ system size is insufficient for error correction of the dynamical qubit, the stabilizers extracted from the measurements are sufficient to detect any single-qubit error. Within the framework of error detection, an operation is considered FT if any single-qubit fault produces a non-trivial syndrome and is thereby detectable [58].
@@
We implement the FBS code on a superconducting quantum processor comprising a $6\times 11$ array of Transmon qubits [59], each coupled to its four nearest neighbors through tunable couplers [60]. The static and dynamical logical qubits are encoded on a 21-qubit subset of the processor, consisting of $N_{d}=9$ data qubits and $N_{a}=12$ ancilla qubits. The data qubits, labeled as $D_{i}$ ($i=1$ to 9), form a $3\times 3$ array and are interlaced with $X$ and $Z$ ancilla qubits for stabilizer measurements. Additional ancilla qubits are incorporated to optimize the QEC circuits. High-fidelity quantum gates and quantum non-demolition (QND) readout [61] are implemented on the 21-qubit subset, with the cumulative distributions of errors illustrated in Fig. 1(c). The experimental setup and device parameters are detailed in the Supplementary Material [62].
Figure 3: Logical gates on the dynamical qubit. (a) FT Pauli gates on the dynamical qubit. Transversal Pauli gates are applied to the data qubits and inserted between stabilizers $S_{Z}^{\rm B}$ and $S_{X}^{\rm C}$, as shown at the top. The final state is characterized using LQST. The state infidelities with error detection after applying the Pauli gates on different encoded states are shown below the circuit. For each dynamical qubit state, the infidelity is averaged over 4 static qubit states {$|0\rangle$, $|1\rangle$, |+ket|$\rangle$| + ⟩, |ket|$\rangle$| - ⟩}. Error bars correspond to 95% confidence intervals. (b) Logical rotation gates on the dynamical qubit around the $Z_{\rm d}$-axis (left) and $X_{\rm d}$-axis (right). The measurement expectation values of the operators $X_{\rm d}$ (cyan), $Y_{\rm d}$ (green) and $Z_{\rm d}$ (orange) are shown as functions of rotation angles $\varphi$ and $\theta$. The gate circuits (shown at the top) for rotations around the $Z_{\rm d}$-axis ($X_{\rm d}$-axis) are inserted between stabilizers $S_{Z}^{\rm B}$ ($S_{X}^{\rm C}$) and $S_{X}^{\rm C}$ ($S_{Z}^{\rm D}$). The filled (open) symbols represent data with (without) error detection in the four stabilizer rounds and the final logical measurement. The dashed curves correspond to trigonometric fits.
@@
We perform period-4 stabilizer measurements to preserve the encoded two-qubit logical states. In each stabilizer round, 4 weight-2 Pauli measurements are conducted. Each measurement result is compared to the result of the same stabilizer type from four rounds earlier, with the first four rounds compared to the initial encoding values. The preservation results for the encoded states $|-,0\rangle$ and $|-,1\rangle$ are presented in Fig. 2(e), showing the measured expectation values of the operator $X_{\rm s}Z_{\rm d}$. With error detection, the logical error rates per round ($\varepsilon_{\rm exp}$), extracted by fitting the exponential decay of the expectation values, are reduced from 14.7% and 14.3% to 2.6% and 2.5%, respectively. The error rates are compared to those of physical two-qubit states (dashed grey lines), with $\overline{\varepsilon}_{\rm phy}=1-\exp(-\tau_{\rm exp}/\overline{T_{2}^{\rm echo}})/2-\exp(-\tau_{\rm exp}/\overline{T_{1}})/2=4.4\%$. Here, $\tau_{\rm exp}=920$ ns is the duration of a stabilizer round, $\overline{T_{2}^{\rm echo}}=11.7~{}\mu s$ is the average spin-echo dephasing time, and $\overline{T_{1}}=77.1~{}\mu s$ is the average energy relaxation time of the physical qubits. Numerical simulation results based on monitored gate and readout errors are also shown for comparison [65]. The larger error rates observed in experiments may arise from leakage-induced correlated errors [66, 67, 68, 69, 70, 71, 64]. We measure an average leakage rate of 0.18% per round on the physical qubits using three-state readout at the end of the circuit [62]. In both the experimental and simulation data, round-dependent oscillations are observed, with higher expectation values occur every four rounds, corresponding to the completion of a full error detection cycle. Beyond error detection, error correction can be applied to the static qubit alone [62]. However, the effectiveness of error correction diminishes for the two-qubit states due to the small code distance of the dynamical qubit and the correlation between errors in the two logical qubits.